Mandy’s Market sells the best coffee, latte, and cappuccino in town while providing the Baristas with a livable wage and health insurance. Their customers love the combination of high quality products and lifestyle issues which both lead to a 90% yearly retention rate. Their customer base averages 4 visits per week and averages spending of $5.25 per visit. The contribution margin is 22%. (Hint: all quantities must be annually)

1. What is the average yearly revenue per customer?

2. Given the retention rate, what is the churn?

3. Based on the churn, what is the Lifespan (BLC)?

4. What is the average repurchase frequency?
Answer #1

1.

Average yearly revenue per customer = 5.25*4*52 = $1092

2.

Churn = 1 - retention rate = 1-90% = 10%

3.

BLC = 1/10% = 10 years

4.

average repurchase frequency = 4 purchases per week